OpenStudy (koolkid98):

Solve the equation for y. 5x – 7y = 10 A. y=-7/5x-7/10 B.y=7/5x+10/5 C.y=5/7x-10/7 D.y=-5x-7 i think D

563blackghost (563blackghost):

We need to isolate y.....so lets write the problem down... \(\huge{5x-7y=10}\) Lets first get rid of `5x` so we would subtract `5x` from both sides... \(\huge{(5x-5x)-7y=-5x+10}\) What would that equal?

OpenStudy (koolkid98):

-7y=-5+10

OpenStudy (koolkid98):

opps im mean -7y=-5x+10

563blackghost (563blackghost):

Nice but dont forget the `x` :) Now we have... \(\huge{-7y=-5x+10}\) Now we need `y` to itself so we would divide by `-7` since it is being multiplied with `y`. So.... \(\huge{\frac{-7y}{-7}=\frac {-5}{-7}x + \frac{10}{-7}}\) Now what would that equal?

OpenStudy (koolkid98):

\[\frac{ -7y }{ -7 }\] would = y or 1y

563blackghost (563blackghost):

-7 over -7 cancels and we are left with y :)